OpenStudy (anonymous):

A store mixes Kenyan coffee worth $10 per kilogram and Turkish coffee worth $14 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 204- kilogram mixture.

OpenStudy (mertsj):

Let x = kg of Kenyan coffee 204-x= kg of Turkish coffee 10x+14(204-x)=12(204)

OpenStudy (mertsj):

Solve the equation.

OpenStudy (mertsj):

\[10x+2856-14x=2448\] \[-4x=-408\] x=102 kg Kenyan coffee 204-x=102=kg Turkish coffee